The net sales and the number of employees for aluminum fabricators with similar characteristics are organized into frequency distributions. Both are normally distributed. For the net sales, the mean is $180 million and the standard deviation is $25 million. For the number of employees, the mean is 1,500 and the standard deviation is 120. Clarion Fabricators had sales of $170 million and 1,850 employees.

  1. Convert Clarion's sales and number of employees to z values. (Negative amounts should be indicated by a minus sign. Round your answers to 2 decimal places.)

Net sales _____

Employees _______

  1. Interpret the z values obtained in part a. (Round your answers to 2 decimal places.)

Net sales are _____ standard deviations _________ the mean

employees are ______ standard deviations _______ the mean

  1. Compare Clarion's sales and number of employees with those of the other fabricators. (Round your answers to 2 decimal places.)

___________ % of the aluminum fabricators have greater net sales compared with Clarion. Only

____________ % have more employees than Clarion.
Answer #1

Solution:-

(a) Given, for net sales mean = $180 million

and standard deviation = $25 million

Calculating z value for net sale,

z = x – μ / σ,

= (180 - 170) / 25 = 0.4

For employees, mean = 1500

and Standard deviation = 120

z = x – μ / σ,

= (1500 - 1850) / 150 = -2.33

(b) Interpreting z value

Net sales are 0.4 (34.46%) standard deviations above the mean.

Employees is -2.33 (.98%) standard deviations below the mean.

(c) 34.46% of aluminium fabricators have greater net sales compared with Clarion.

0.98% have more employees than Clarion.